Sex and violence TheJournal of Applied Psychology reported on a study that examined whether the content of TV shows influenceYou are a consultant to the marketing department of a business preparing to launch an ad campaign for a new product. The company can afford to run ads during one TV show, and has decided not to sponsor a show with sexual content. Do these results indicate that viewer memory for ads may differ depending on program content? Conduct an appropriate hypothesis test at the a= 0.05 level. Show your working, which test statistic are u using? Assume population standard deviations are equal. What would you recommend?
Answer #1

Since we need to compare more than three population mean so one way ANOVA will be used.

Hypotheses are:

H_{0}:\mu_{1}=\mu_{2}=\mu_{3}

Ha: At least one population mean differ.

Here we have

108, T1 2.08 +1.713.17 2.08, 72 - 1.71,73 3.17,si 1.87, s 1.762, s2 = 1.772, 2.32

Now :

SStreatments _ n! (zi-i)2 + n2(T2-x)2 + n3Tz-x)2-124.4376

and

SSerror = (n1-1) sỉ + (n2 - + (R3-1)s-1040.8318 1) s

Therefore

SStotal-SStreatments SSerror-1165.2694

Since there are 3 different groups so we have k=3. Therefore degree of freedoms are:

dftreat ments-k-1 = 3-1 = 2

dferrorN - k-324- 3321

dftotal - 323

-------------

Now

-622188 sst re at ments MStreatmentsdftreatments

Serror = error = 3.2425 MSerror dferror eTTo eTTo

F test statistics is

Streatments19.1887 M Serror

So p-value of the test is 0.0000. Since P-value is less than 0.05 so we reject the null hypothesis. That is we can conclude that viewer memory for ads may differ depending on program content.